# A Generalized Theory for Optical Cooling of a Trapped Atom with Spin

### v2

@@ V Analysis
The formalism in this paper reveals that all two-photon cooling schemes share a fundamental similarity: they involve a Raman transition between neighboring motional states with different spin states. Close-to-resonance cooling schemes like PG and GM utilize light shifts to bring the states into resonance, while further detuned cooling methods, such as Raman sideband (RSB) cooling, use the two-photon detuning of two frequencies and two far detuned ground-states. Similarly, techniques like $\Lambda$-GM cool by establishing dark states states through destructive interference of two frequencies and two far detuned ground-states.
@@
PG cooling is fundamentally limited to the optical pumping asymmetry, which for $F$ to $F+1$ is $\langle n\rangle\approx 1$. GM and $\Lambda$-GM lower the population to $\langle n\rangle\approx\eta^{2}$. But, spontaneous emission in these schemes still drives the population out of the dark state through either spontaneous emission from the first-order LD parameter or in the effective ground-state formalism, the position dependent component of the collapse operator. The dark $n=0$ state is thus not completely dark, as it still can be driven through the $n=1$ excited state. However, when using the Raman light in an optical pump, this is unpreventable, as the light requires a position-dependent term to produce the Raman coupling between motional states.
@@
RSB cooling represents the ultimate limit of cooling, detuning so far from the excited state that the spontaneous emission of the Raman coupling light is negligible. The re-pumping is achieved through a separate spontaneous emission beam optimized to make the dark state truly dark.
@@
In practical laboratory settings, PG, GM, and $\Lambda$-GM are more convenient due to their operational simplicity and similarity to the MOT configuration. But RSB is still employed for high ground-state preparation despite its significant experimental overhead. This paper demonstrates that Grey molasses cooling combined with the $\Lambda$-enhancement can achieve significant ground-state populations. This raises the question of how schemes resembling MOT configurations can be adapted for near-ground-state cooling. Likely, this adaptation involves adding an additional beam specialized for optical pumping. For example adding an additional circularly polarized beam close-to-resonance beam for optical pumping (not counter-propagating). This should, from this model, provide the optical pumping without driving sidebands of the dark state. Then in the limit of large detuning for the GM beams, it would effectively create a $L_{+}=\sqrt{\gamma_{+}}e^{ik\hat{x}}\hat{F}_{+}$ from the RSB model. This will be the investigation of future experimental and theoretical work.
@@ VI Conclusion
This study presents a comprehensive theoretical framework for major cooling mechanisms in neutral atom tweezers, uncovering shared principles across diverse techniques. Our approach combines detailed full-level structure simulations with a simplified spin model, offering novel insights for optimizing cooling schemes. By extending previous research to encompass arbitrary spins and polarizations, we show that gray molasses cooling may potentially reach the ground-state, challenging conventional limits and opening up new avenues for exploration.
Future research will investigate polarization gradients in three-dimensional beam configurations and examine deviations from the idealized one-dimensional cases presented here. We also plan to conduct simulations for specific alkali atoms, including all relevant ground and excited states, to assess the impact of multiple excited states and cross-coupling between lasers addressing different hyperfine ground-states. These simulations will help establish fundamental cooling limits for each atomic species for various techniques. Future research will explore innovative cooling approaches that blur the boundaries between established techniques.
Rapid, robust, and efficient ground-state optical cooling will increase the coherence and fidelity of experiments throughout the ultracold field. Reducing thermal motion in Rydberg atoms will pave the way for higher-fidelity quantum gates. Rapid optical cooling techniques could also reduce the long times associated with thermal evaporation for producing degenerate gases.
@@ Acknowledgements
which are on the order of unity and give the strength of a $F_{g}$ to $F_{e}$ transition relative to the $J_{g}$ to $J_{e}$ transition with $\sum_{F_{e}}|O^{J_{g}J_{e}}_{IF_{g}F_{e}}|^{2}=1$
@@
$C^{(0)}=(-1)^{3F-F^{\prime}}\sqrt{\frac{1}{3}}\frac{2F^{\prime}+1}{\sqrt{2F+1}}$ (92)
@@
$C^{(2)}=(-1)^{3F-F^{\prime}}\sqrt{\frac{30(2F^{\prime}+1)}{F(F+1)(2F+1)(2F-1)(2F+3)}}$ (94)
×$\times\begin{pmatrix}F&1&F^{\prime}\\ 2&F&0\end{pmatrix}\left|\begin{Bmatrix}1&F&F^{\prime}\\ 0&1/2&F\end{Bmatrix}\right|^{2}.$

### v3

@@ V Analysis
In this paper, we have demonstrated that polarization gradient (PG), gray molasses (GM), $\Lambda$-enhanced gray molasses ($\Lambda$-GM), electromagnetically induced transparency (EIT), and Raman sideband (RSB) cooling can all be understood within a unified framework of spin cooling. Our formalism reveals that all these two-photon cooling schemes share a fundamental similarity: they involve a Raman transition between neighboring motional states with different spin states.
After adiabatically eliminating the excited state and expanding to first order in the Lamb-Dicke (LD) parameter, each cooling model exhibits a similar structure, featuring a Hamiltonian and collapse operator with both position-independent and position-dependent terms. The key ingredients for effective cooling are: (1) a differential light shift between spin states equal to the trapping frequency, ensuring resonance between $|g_{2},n\rangle$ and $|g_{1},n-1\rangle$, (2) preferential optical pumping ($\gamma_{-}/\gamma_{+}<1$) to the lower energy state, and (3) coherent coupling between the resonant states.
We have identified three primary factors that determine the final temperature: a large asymmetry in optical pumping, a small LD parameter, and maximizing the darkness of the lower energy state. Different cooling techniques implement these principles in various ways, with close-to-resonance cooling schemes like PG and GM utilizing light shifts to bring the states into resonance, while further detuned methods such as RSB cooling use the two-photon detuning of two frequencies and two far-detuned ground states.
PG and GM are experimentally efficient, using a pair of counter-propagating beams for all three requirements. However, PG cooling is fundamentally limited by unity pumping asymmetry, with $\langle n\rangle\approx\gamma_{-}/\gamma_{+}\approx 1$ for $F$ to $F+1$ transitions. GM differs in the darkness of the lower energy state, utilizing natural dark states of the ground state manifold when $F^{\prime}\leq F$. In the limit of $\gamma_{-}\ll\gamma_{+}$, GM’s fundamental limit is $\langle n\rangle\approx\eta^{2}\gamma_{h}/\gamma_{+}$, where $\gamma_{h}/\gamma_{+}$ is the ratio of the effective scattering rate of the dark state to the pumping rate, typically close to unity for GM. For heavy atoms with $\eta<0.1$, this can lead to ground state preparation.
$\Lambda$-GM maintains the same scaling as GM but enhances cooling by introducing more dark states and a second coherent laser to address the other hyperfine ground state. The relative detuning in these EIT configurations in principle relaxes the condition that light shifts must equal the trapping frequency (the case for GM), as the relative detuning of the two lasers can compensate.
@@
However, these schemes still rely on these same two counter-propagating lasers for optical pumping. Despite the improvements offered by GM and $\Lambda$-GM, spontaneous emission in these schemes still drives the population out of the dark state through either spontaneous emission from the first-order LD parameter or, in the effective ground state formalism, the position-dependent component of the collapse operator. The dark $n=0$ state is thus not completely dark, as it can still be driven through the $n=1$ excited state.
@@
RSB cooling represents the ultimate limit of cooling, detuning so far from the excited state that the spontaneous emission of the Raman coupling light is negligible. It separates the cooling process into three distinct steps: (1) the relative detuning of two beams brings the trapping frequencies of two spin states into resonance, (2) these beams create coherent transfer, and (3) an additional resonant laser provides optical pumping, specifically designed to maximize the darkness of the dark state. The re-pumping is achieved through a separate spontaneous emission beam optimized to make the dark state truly dark.
@@
In practical laboratory settings, PG, GM, and $\Lambda$-GM are more convenient due to their operational simplicity and similarity to the magneto-optical trap (MOT) configuration. However, RSB is still employed for high ground state preparation despite its significant experimental overhead. Our work demonstrates that gray molasses cooling combined with the $\Lambda$-enhancement can achieve significant ground state populations, raising the question of how schemes resembling MOT configurations can be adapted for near-ground state cooling.
One potential strategy would be to use GM or $\Lambda$-GM, detuned far enough to make the scattering rate negligible, then add a resonant beam for spontaneous emission. The goal is to eliminate the position-dependent collapse operator that drives the dark transition, which is the only coupling out of the dark ground state. This perhaps could be accomplished by having separate optical pumping beams that are co-propagating and drive only the bright state. In the simple model picture, this would effectively create a collapse operator $L_{+}=\sqrt{\gamma_{+}}e^{ik\hat{x}}\hat{F}_{+}$ similar to the RSB model, where the second term doesn’t drive the dark state.
This approach could potentially combine the advantages of both worlds, offering the experimental simplicity of GM and $\Lambda$-GM with the improved ground state preparation of RSB. Further investigation of this possibility will be the subject of future experimental and theoretical work, as we continue to explore how these cooling schemes can be optimized and combined to achieve even lower temperatures and higher ground state populations in trapped atomic systems.
@@ VI Conclusion
This study presents a comprehensive theoretical framework for major cooling mechanisms in neutral atom tweezers, uncovering shared principles across diverse techniques. Our approach combines detailed full-level structure simulations with a simplified spin model, offering novel insights for optimizing cooling schemes. By extending previous research to encompass arbitrary spins and polarizations, we show that gray molasses cooling may potentially reach the ground state, challenging conventional limits and opening up new avenues for exploration.
Future research will investigate polarization gradients in three-dimensional beam configurations and examine deviations from the idealized one-dimensional cases presented here. We also plan to conduct simulations for specific alkali atoms, including all relevant ground and excited states, to assess the impact of multiple excited states and cross-coupling between lasers addressing different hyperfine ground states. These simulations will help establish fundamental cooling limits for each atomic species for various techniques. Future research will explore innovative cooling approaches that blur the boundaries between established techniques.
Rapid, robust, and efficient ground state optical cooling will increase the coherence and fidelity of experiments throughout the ultracold field. Reducing thermal motion in Rydberg atoms will pave the way for higher-fidelity quantum gates. Rapid optical cooling techniques could also reduce the long times associated with thermal evaporation for producing degenerate gases.
@@ Acknowledgements
which are on the order of unity and give the strength of a $F_{g}$ to $F_{e}$ transition relative to the $J_{g}$ to $J_{e}$ transition with $\sum_{F_{e}}|O^{J_{g}J_{e}}_{IF_{g}F_{e}}|^{2}=1$. For more details, see Refs. [40] and [69].
@@
$C^{(0)}=(-1)^{3F-F^{\prime}+1}\sqrt{\frac{1}{3}}\frac{2F^{\prime}+1}{\sqrt{2F+1}}$ (91)
@@
$C^{(2)}=(-1)^{3F-F^{\prime}}\sqrt{\frac{30}{F(F+1)(2F+1)(2F-1)(2F+3)}}$ (93)
×$\times(2F^{\prime}+1)\times\begin{pmatrix}F&1&F^{\prime}\\ 1&F&2\end{pmatrix}\left|\begin{Bmatrix}1&F&F^{\prime}\\ 0&1/2&F\end{Bmatrix}\right|^{2}.$